OpenStudy (anonymous):

A cyclist is acceleratingat 5m/s in 30 seconds. The cyclist initial speed is .50m/s. What will be the cyclist final speed?

OpenStudy (anonymous):

v(final) = v(initial) + at v(final)= .50 m/s + (5)(30) =

OpenStudy (anonymous):

Thats easy, the final speed is 200 m/s Fs=Is+at Fs=50 m/s + 5m/s^2 (30 s) Fs=200 m/s

OpenStudy (mrnood):

that is not correct if the question is correctly written. It says the change in speed is 5m/s (and it takes 30 s to achieve that) so the final speed is 5.5 m/s (IF it was meant to be 5 m/s^2 for 30 s then the above answers are correct - but it would have to be a rocket powered bike to maintain that acceleration ) @Yocy - you have misread it anyhow - the initial speed is 0.5 m/s (not 50 m/s)

OpenStudy (mrnood):

BTW - if he DID accelerate at 5 m/s^2 for 30 s then he would be doing 150.5 m/s after 30s - which is 336mph So I think the question IS set correctly and the total change in speed is 5m/s

OpenStudy (anonymous):

Oh yeah you're right MrNood ^^ I didn't see the period before the 50. Sorry